# Estimating energy levels from lattice QCD correlation functions using a transfer matrix formalism

### v1

@@ II.1 Procedure for spurious eigenvalue filtration
    First, we perform bootstrap resampling on the original dataset to generate $N_{b}$ bootstrap samples. For each of these samples, we carry out GEVP calculations, as defined in Eq. (9), and compute the energy eigenvalues $\lambda_{n}^{m,b}$. Here, the index $n$ enumerates the eigenvalues obtained for the $m$-th iteration and the $b$-th bootstrap sample. The ranges are defined as $0\leq n\leq m$, $1\leq m\leq\lfloor(n_{T}+1)/2\rfloor$, and $b\in\{1,2,\dots,N_{b}\}$.
@@
    Next, we eliminate all eigenvalues with $\left|\arg\lambda_{n}^{m,b}\right|>\epsilon$, retaining only the real eigenvalues (value of $\epsilon$ depends on the required precision). Eigenvalues outside the range $0<\lambda\leq 1$ are discarded, as they do not correspond to physically meaningful energy states. For staggered valence quarks, however, because of the oscillating states, one needs to include eigenvalues in the range $\lambda_{n}^{m,b}\in(-1,0)$.
@@
    For each iteration, a histogram is generated from the eigenvalues obtained across all bootstrap samples. To reduce statistical fluctuations in the histogram, smoothing techniques such as Gaussian filtering are utilized. The peaks in the histogram are then identified by determining their centers and calculating their full width at half maximum (FWHM). These peak-centers and their corresponding widths are denoted by $C_{r}^{m}$ and $\Delta_{r}^{m}$, respectively, where $m$ refers to the iteration number, and the subscript $r$ indexes the peaks in descending order of their center values. This approach allows for a clearer identification of dominant eigenvalues, An example plot demonstrating this process is provided in Appendix D.
@@
    Each peak thus identified in the histogram represents an eigenstate of the transfer matrix. The peak-centers and their widths are then utilized to guide the pivoting procedure for selecting eigenvalues, ensuring that only physically meaningful eigenstates are retained. This procedure enhances the reliability of the extracted energy levels by systematically filtering out spurious eigenvalues and maintaining a consistent ordering of eigenvalues.
@@
    For each peak in the $m$-th iteration, we select the eigenvalue $\lambda_{n}^{m,b}$ that is closest to $C_{r}^{m}$ and lies within the range $\left[C_{r}^{m}-\Delta_{r}^{m},C_{r}^{m}+\Delta_{r}^{m}\right]$ for each bootstrap sample. This procedure results in a set of filtered eigenvalues, denoted as $\tilde{\lambda}_{r}^{m,b}$, where the index $r$ represents the peaks in the histogram, or equivalently, the well-resolved states in the spectrum of $T^{m}$. For instance, $-\log\tilde{\lambda}_{r=0}^{m,b}$ represents the ground-state energy eigenvalue for the $m$-th iteration and the $b$-th bootstrap sample.
@@
Naturally, a concern here is on the accurate estimation of errors in the eigenvalues obtained using the above procedure, particularly when comparing them with effective masses. One needs to account for the potential influence of outliers in the bootstrap eigenvalue estimates. In Refs. [1, 19] a double-bootstrap method was utilized to address that. Instead, we adopt the 68th percentile as the bootstrap confidence interval. Despite the filtration procedure applied to the eigenvalues $\tilde{\lambda}_{r}^{m,b}$, some outliers may still remain. These outliers can skew the bootstrap mean and inflate the bootstrap standard error, particularly in certain iterations. To mitigate this, we define the bootstrap confidence interval by using the 68th percentile, which corresponds to the interval between the 16th and 84th percentiles of the bootstrap distribution. The mean of the eigenvalues within this 68th percentile interval is then quoted as the bootstrap mean. However, to be conservative with errors, we estimate those using the 68th percentile of the full samples without removing any outlier. We find this approach reduces the influence of extreme values robustly with reliable errorbars, leading to more stable and robust estimates. We automate the whole procedure described above after fixing the count-threshold for peak detection and the smoothing-parameter of the histogram counts, thus reducing subjectiveness and associated systematics.
III Results
We now present the results as obtained by the aforementioned TGEVP method. Eigenvalues are computed using two different approaches, namely by PYTHON [25] and FEAST eigensolvers [26], ensuring cross-checking while maintaining high precision evaluation. In Figs. [1, 2, 3, 4] we show our results. In each of the figures we present the lowest eigenvalue as estimated by the above method, along with the standard effective mass, defined as $m_{eff}(t)={\mathrm{log}}[C(t)/C(t+1)]$. In each case, throughout this article, blue circles represent the regular effective mass $m_{eff}(t)$, and the red circles represent the lowest eigenvalues as computed over the iterations. The magenta band represents the correlated fit of the two-point function using a single exponential, illustrating the fit range and the associated $1\sigma$ error. The faded-green band corresponds to the eigenvalues, fitted with a constant term over multiple iterations, also using a correlated fit. This comparison highlights the stability and error behavior of the TGEVP eigenvalues over regular effective mass approach. A large number of other results, including the first excited states, are shown in Appendix C and are discussed there.
III.1 Case I: Charmonia
We first present the results for the 1S-charmonia in Fig. 1. The results corresponds to the two-point correlation functions of $0^{-}$ (top) and $1^{-}$ (bottom) states of charmonia, as computed using overlap valence quarks with a HISQ lattice ensemble ($48^{3}\times 144$, $a=0.0582$ fm, $n_{meas}=180$). As can be observed, while the effective mass is fluctuating over time-slices, the lowest eigenvalue estimated over different iterations is very stable. In Appendix C, we further show similar results for other lattice ensembles along with the first excited states. We also show the continuum extrapolation of mass splittings between the excited and the ground states, and compare those with their experimental values.
Figure 1: Comparison of the results between the standard effective mass of the ground state (blue circles) and the lowest eigenvalue (red circles) corresponding to the two-point correlation functions of the pseudoscalar ($J^{P}\equiv 0^{-}$) (top) and vector ($1^{-}$) (bottom) states of charmonia. Number of measurments ($n_{meas}$) in this case is $180$.
III.2 Case II: Heavy Baryons
In Fig. 2, we show similar results obtained from two-point correlation functions of the heavy baryon $\Xi_{cc}$, at two different quark masses. It again very clearly shows the stability of the lowest eigenvalue over the iteration numbers, as opposed to unstable and fluctuating effective mass data. In such cases, the TGEVP results can really be helpful to extract the ground state, avoiding systematic uncertainty due to various choices of fit ranges in the regular fitting procedures.
Figure 2: Similar to Fig. 1, corresponding to smeared-smeared two-point functions for the doubly charmed baryon $\Xi_{cc}$, computed at two different light quark masses with number of measurements ($n_{meas}$) $1612$ and $2000$ respectively.
III.3 Case III: Nucleon
In Fig. 3, we show similar results obtained from a smeared-point two-point function of the nucleon computed at a near physical pion mass. While the results from the two estimations are consistent, the eigenvalues at the larger iterations have smaller errors compared to errors in the effective mass at the larger times.
Figure 3: Similar to Fig. 1, corresponding to a smeared-point two-point function of nucleon $(n_{meas}=300)$
III.4 Case IV: Nuclei
One of the main bottlenecks in lattice QCD studies of nuclei is the lack of reliability in extracting the energy levels, even for the ground states of low-lying nuclei [12, 13, 14, 15]. This is primarily due to poor signal-to-noise ratio in the two-point correlation functions at lighter quark masses, resulting very noisy and small plateau at a large distance in the effective mass plots. As a result, statistical error together with fit-uncertainties, becomes large making the whole process unreliable. It is thus quite important to check if TGEVP method can be helpful in this. In Fig. 4, we show the effective mass of the ground state and lowest eigenvalue of TGEVP spectrum for a few nuclei. These results are for 2H (top), 3He (middle) and 4He (bottom) (more such results are shown in Appendix C). The two-point functions are obtained with overlap valence quarks on the $N_{f}=2+1+1$ HISQ gauge ensembles ($0.058248^{3}\times 144,~{}a=$ 0.0582 fm) at the quark masses, $m_{u}=m_{d}=m_{s}$. The method for computing the correlation functions for these nuclei has been provided in Ref. [15]. As can be observed that the effective masses of these nuclei exhibit significant noise from the start, requiring substantial systematic error corrections in traditional fitting methods. In contrast, the TGEVP eigenvalues remain stable across numerous iterations. With a large number of measurements, it is expected that the erorrs in eigenvalues in each iteration can be reduced. However, it is also necessary to assess the method’s effectiveness for the cases with closely spaced states carrying finite momentum, particularly when dealing with varied source-sink setups in nuclear systems. Another point of concern is correlation over iterations which we will discuss next. Nevertheless, as the results suggest, this method holds excellent promise for extracting reliable energy levels for nuclei in the future.
III.5 Correlation in TGEVP eigenvalues
A natural concern is for accounting the correlation between eigenvalues obtained across iterations and the validity of a constant fit over iterations. Addressing this correlation mathematically, where data in each step in a matrix are generated randomly from the previous step, is non-trivial. In Refs. [1, 19] a double-bootstraping is done to address that. In this work, we define the bootstrap confidence interval by using the $68$th percentile. To assess the extent of correlation, we define the correlation matrix under bootstrap and compute that for the eigenvalues that lies within the 68th percentile interval. In Fig. 5, we show such a color-coded correlation-plot for the two-point correlation functions for $\eta_{c}$ that was shown in the top pane of Fig. 1. As observed, the corelation of eigenvalues across iterations is insignificant at the beginning and increases towards the end. Particularly in the fit region (Fig. 1), it remains small enough to justify a correlated fitting with a constant term. Additional examples are provided in the Appendix E, showing smaller correlations for most datasets, except for nuclei where relatively larger correlations are observed. For nuclei, however, while this method does not solve the SNR problem, it mitigates that to an extent and provides a more reliable estimation of the energy levels compared to existing fitting methods.
Figure 5: Color-coded correlation matrix showing the correlations between the eigenvalues across iterations. This dateset corresponds to the two-point correlation functions of the lowest $0^{-}$ state ($\eta_{c}$ meson), showed in the top pane of Fig. 1.
IV Conclusion and outlook
This work presents an efficient method for extracting energy levels by applying generalized eigenvalue framework to the transfer matrix constructed from time-shifted two-point correlation functions of lattice QCD. We find this method is equivalent, both analytically as well as numerically, to the Lanczos method [1], but is more general and easy to implement. A pivoting procedure, along with a bootstrap confidence interval of 68th percentile, is employed for removing spurious eigenvalues, and is found to be more efficient than the Cullum-Willoughby method. We demonstrate the efficacy of the method by calculating first two eigenvalues from two-point functions of a variety of hadrons, including nuclei. For charmonia, along with the individual energy levels we also compute energy splittings between different states and show that, when extrapolated to the continuum limit, they align well with experimental values, even with limited statistics. In most cases with higher statistics, we find the correlation between eigenvalues across iterations is sufficiently low, allowing for a constant fit, which yields more reliable results than traditional exponential fitting methods. For nuclei, we find the correlation between eigenvalues across iterations to be higher than that for single nucleons, particularly when statistics are low. The increased correlation poses a challenge for solving the SNR problem, and a thorough evaluation is necessary to account this correlation for accurate estimation of energy levels in such cases. However, within the limit of a given statistics, the TGEVP eigenvalues remain more stable across iterations compared to the noisy effective masses obtained through traditional methods. This stability suggests that with increased statistics, the use of a multi-operator setup, and proper accounting of correlations, this procedure holds promise for accurately determining energy levels even in challenging multi-nucleon systems.
Acknowledgment
This work is supported by the Department of Atomic Energy, Government of India, under Project Identification Number RTI 4002. Computations related to some of the correlators were done on the Cray-XC30 of ILGTI, TIFR (which has recently been closed), and the computing clusters at the Department of Theoretical Physics, TIFR, Mumbai. NM is thankful to Keh-Fei Liu, Gunnar Bali and Wolfgang Soeldner for providing us with the nucleon, and heavy baryon data. NM and DC also acknowledge discussions with Piyush Srivastava at various stages of this work. We would also like to thank Ajay Salve, Kapil Ghadiali, and T. Chandramohan for computational support.
Appendix A Equivalence between Lanczos and TGEVP methods
Assuming $|u_{i}\rangle$, $i=0,1,\cdots,m$, form a complete set of orthonormal basis in the Krylov-subspace $\mathcal{K}_{m}$, the vectors $|v_{i}\rangle$, that enter in Eqs. (7) and (8), can be expanded as,
$|v_{i}\rangle=\sum_{j=0}^{m}\alpha_{ij}|u_{j}\rangle.$ (10)
where $\text{where }\tilde{T}_{ij}=\langle u_{i}|\mathcal{T}|u_{i}\rangle.$
We need to show that the above generalized eigenvalue problem is exactly equivalent to the eigenvalue problem in Eq. (A). Starting with (9), this equivalence can be demonstrated through the following steps,
$T_{ij}(x_{n})_{j}$ $=$ $\lambda_{n}V_{ij}(x_{n})_{j}\,,$ (12)
$\tilde{T}_{il}\alpha_{jl}(x_{n})_{j}$ $=$ $\lambda_{n}\alpha_{ji}(x_{n})_{j}\,,$ (15)
Here, in the second step, we have substituted $|v_{i}\rangle$ utilizing Eq. (10). In the third step, we have multiplied the both sides of the equation using inverse of $\alpha_{ij}$ (given that $|v_{i}\rangle~{}$ are linearly-independent, inverse of the matrix $\alpha_{ij}$ exists.). In the final step, we have redefined $(\tilde{x}_{n})_{i}=\alpha_{ji}(x_{n})_{j}$ to arrive at the desired form.
Appendix B Numerical comparison between Lanczos and TGEVP methods

### v2

@@ III.2 Adaptive Kernel density Estimator
    Pilot Estimation: We begin by constructing a pilot estimate using standard KDE with a globally optimized bandwidth $h_{\text{global}}$. The global bandwidth, $h_{\text{global}}$, is chosen using Silverman’s rule of thumb [36, 34, 38], which is given by $h=0.9\times\min\left(\sigma,\frac{\text{IQR}}{1.34}\right)\times n^{-1/5}$, where IQR is the inter-quartile range, $\sigma$ is the standard deviation and $n$ is the sample size. For the adaptive KDE procedure, we focus on using this to generate the pilot estimate for each data point.
@@
    Bandwidth Modulation: After obtaining the pilot estimate, the local bandwidth, $h_{i}$, for each point, $x_{i}$, is calculated adaptively based on the density of points in the vicinity of $x_{i}$ as,
    $h_{i}=h_{\text{global}}\left(\frac{f_{\text{pilot}}(x_{i})}{g}\right)^{-\alpha},$ (15)
    where $f_{\text{pilot}}(x_{i})$ is the pilot density at the point $x_{i}$, and $g$ is the geometric mean of the pilot densities across all data points:
    $g=\exp\left(\frac{1}{n}\sum_{i=1}^{n}\log f_{\text{pilot}}(x_{i})\right).$ (16)
    The parameter $\alpha$ adjusts the strength of the adaptivity.
@@
    Adaptive KDE Estimation: The final adaptive KDE is then computed using the locally adjusted bandwidths. The density estimate at each point $x$ is:
    $\hat{f}(x)=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{h_{i}}K\left(\frac{x-x_{i}}{h_{i}}\right),$ (17)
    where $K$ is the kernel function, which we choose to be Gaussian for this work. This step incorporates the locally adaptive bandwidths, allowing for finer resolution near denser regions and smoother estimates in sparser regions.
@@
    Choosing the Optimal $\alpha$: To determine the optimal $\alpha$ in our adaptive KDE framework, we employ the method of $K$-fold cross-validation [36, 34, 38], a widely utilized technique to assess the performance of a model and to mitigate over-fitting. In $K$-fold cross-validation, the entire dataset is partitioned into $K$ almost equal-sized, mutually-exclusive datasets. For each datasets $k\in\{1,\cdots,K\}$, we perform the following procedure:
    1. (a)
      Treat the $k$-th dataset as the validation set and rest of the $(K-1)$ dataset as the training set.
    2. (b)
      Using the training set, we estimate the density $\hat{f}(x;\alpha)$ with a given $\alpha$ via our adaptive KDE procedure which is laid out in the first three steps.
    3. (c)
      Then, the robustness of the density estimated by the training set is evaluated by computing negative log-likelihood of the validation set as below,
      $\mathcal{L}^{k}(\alpha)=-\sum_{x_{i}\in\text{validation set}}\log\hat{f}(x_{i};\alpha).$ (18)
    4. (d)
      This procedure is repeated for each of the $K$ data partitions, and the average cross-validated log-likelihood is estimated as,
      $CV(\alpha)=\frac{1}{K}\sum_{k=1}^{K}\mathcal{L}^{k}(\alpha).$ (19)
    We repeat the above procedure for a large number of $\alpha$ values on a grid. The optimal $\alpha^{\ast}$ is then chosen as the one that minimizes the average $CV(\alpha)$. Finally this optimal $\alpha^{\ast}$ is plugged into Eq.(17) to perform final density estimation using adaptive KDE. Note that the entire procedure, including the choice of $\alpha$, as outlined above is data-driven and does not need any manual tuning, making this method robust and the preferred choice.
This adaptive construction provides finer resolution in regions where the eigenvalue distribution is sharply peaked, thereby capturing closely spaced modes, while applying greater smoothing in the sparse tails, where statistical noise dominates. In effect, it allows the KDE to adapt to the spectral structure revealed by the bootstrap ensemble, offering a statistically robust and automated tool for the identification and rejection of spurious modes.
As with histogram-based approaches, we emphasize that the role of KDE in our workflow is not merely visualization, but rather to enable quantitative detection of peak locations and widths. These features directly inform the thresholds used in our filtering strategy, ensuring that our treatment of spurious modes is grounded in the underlying statistical structure of the spectrum.
III.3 Analysis of eigenvalue distributions with the Histogram and KDE methods
In this subsection, we apply the previously described histogram-based techniques and the KDE method to analyze the eigenvalue distributions and assess their effectiveness in filtering out spurious eigenvalues produced by the TGEVP method. We choose two examples: two-point correlation functions of nucleon and $J/\psi$ meson. The goal is to demonstrate how these methods can be utilized to estimate the FWHM of the peaks in the distributions and remove outliers based on these estimates.
We begin by constructing a histogram of the eigenvalues, which are shown in Fig. 1. The histograms reveal the distribution of eigenvalues, with some statistical noise due to the discrete nature of the data. Following the procedure described in III.1, we then apply a Gaussian smoothing filtration to the histograms. The smoothed distribution is shown by the blue line, highlighting the central features of the distribution while suppressing fluctuations.
Next we compute the kernel density estimate, as described in III.2, and the resulting distribution is depicted by the red line alongside the smoothed histograms in Fig. 1. We find the adaptive KDE result provides a continuous estimate of the probability density function of the eigenvalues and captures the shape of the distribution more accurately than the discrete histogram. Note the adaptive nature of the KDE adjusts the bandwidth depending on the local density of the data, yielding a smoother and more reliable density estimate. To ensure an optimal balance between smoothness and fidelity, we determine the regularization parameter, $\alpha$, using the $K$-fold cross-validation method, as described in the previous subsection. In Figure 2, we present the results of this optimization procedure, with optimal values of $\alpha$, for both the cases of nucleon and the $J/\psi$ meson correlators. For all other correlators investigated (as presented in the Results section), we observe similar advantages of the adaptive KDE method, achieved in a fully data-driven manner through hyperparameter optimization. Based on these findings, we recommend the adaptive KDE method as the most robust and reliable approach for eigenvalue filtration. The Gaussian-convoluted method serves as a complementary option, useful for cross-verification.
Figure 1: Histogram, Gaussian-smoothed histogram, and kernel density estimate (KDE) for the eigenvalue distribution of the $1^{-}$ charmonium (top panel) and nucleon (top panel) 2-point correlation function. The histogram shows the bootstrapped eigenvalue distribution, the smoothed histogram suppresses noise, and the KDE provides a continuous estimate of the probability density.
Figure 2: Cross-validation score $\text{CV}(\alpha)$ as a function of the regularization parameter $\alpha$ for $1^{-}$ charmonium (top panel) and single nucleon (bottom panel). The minimum of the curve indicates the optimal value $\alpha^{\star}$. The red marker denotes the location of the minimum cross-validation score.
IV Flowchart for filtration of spurious eigenvalues
Having broadly discussed the methodologies for eigenvalue filtration in the previous section, we now present a complete flowchart outlining the implementation of the above-mentioned procedures for extracting energy levels within our proposed TGEVP framework.
  1. 1.
    First, perform a bootstrap resampling on the original dataset to generate $N_{b}$ bootstrap samples. For each of these samples, one needs to carry out GEVP calculations, as defined in Eq. (8), computing the energy eigenvalues $\lambda_{n}^{m,b}$. Here, the index $n$ enumerates the eigenvalues obtained for the $m$-th iteration and the $b$-th bootstrap sample. The ranges are defined as, $0\leq n\leq m$, $1\leq m\leq\lfloor(n_{T}+1)/2\rfloor$, and $b\in\{1,2,\dots,N_{b}\}$.
  2. 2.
    Next, we eliminate all eigenvalues for which $\left|\arg\lambda_{n}^{(m,b)}\right|>\epsilon$, retaining only those that are real within numerical precision (the value of $\epsilon$ is set based on the required floating-point tolerance). Eigenvalues outside the range $0<\lambda\leq 1$ are discarded, as the transfer matrix is a positive operator. However, if the correlation functions are affected by poor statistics, the removal of negative and complex eigenvalues may introduce statistical bias in the final energy estimates. To guard against this, we provide diagnostic tools for evaluating the impact of such removals, with further details presented in Appendix D.
  3. 3.
    For each iteration, estimate the density of eigenvalues obtained across all bootstrap samples using adaptive KDE III.2 and/or Gaussian-convoluted histogram III.1 method as described in the previous section (note KDE is our recommended method for its robustness). The peaks in the estimated eigenvalue density are then identified by determining their centers and calculating their full width at half maximum. These peak-centers and their corresponding widths are denoted by $C_{r}^{m}$ and $\Delta_{r}^{m}$, respectively, where $m$ refers to the iteration number, and the subscript $r$ indexes the peaks in descending order of their center values.
  4. 4.
@@ IV Flowchart for filtration of spurious eigenvalues
    Each peak thus identified in the estimated eigenvalue density represents an eigenstate of the transfer matrix. The peak-centers and their widths are then utilized to guide the procedure for selecting eigenvalues, ensuring that only physically meaningful eigenstates are retained. This procedure enhances the reliability of the extracted energy levels by systematically filtering out spurious eigenvalues and maintaining a consistent ordering of eigenvalues.
@@
    For each peak in the $m$-th iteration, then select the eigenvalue $\lambda_{n}^{m,b}$ that is closest to $C_{r}^{m}$ and lies within the range $\left[C_{r}^{m}-3\Delta_{r}^{m},C_{r}^{m}+3\Delta_{r}^{m}\right]$ for each bootstrap sample. This procedure results in a set of filtered eigenvalues, denoted as $\tilde{\lambda}_{r}^{m,b}$, where the index $r$ represents the peaks in the histogram, or equivalently, the well-resolved states in the spectrum of $T^{m}$. For example, $-\log\tilde{\lambda}_{r=0}^{m,b}$ represents the ground state energy eigenvalue for the $m$-th iteration and the $b$-th bootstrap sample.
@@
Naturally, a concern here is on the accurate estimation of errors in the eigenvalues obtained using the above procedure, particularly when comparing them with traditional method such as, effective masses. One needs to account for the potential influence of outliers in the bootstrap eigenvalue estimates. In Refs. [1, 25] a double-bootstrap method was utilized to address that. Instead, we adopt the 68th percentile as the bootstrap confidence interval. Despite the filtration procedure applied to the eigenvalues $\tilde{\lambda}_{r}^{m,b}$, some outliers may still remain. These outliers can skew the bootstrap mean and inflate the bootstrap standard error, particularly in certain iterations. To mitigate this, we define the bootstrap confidence interval by using the 68th percentile, which corresponds to the interval between the 16th and 84th percentiles of the bootstrap distribution. The mean of the eigenvalues within this 68th percentile interval is then quoted as the bootstrap mean. However, to be conservative with errors, we estimate those using the 68th percentile of the full samples without removing any outlier. We find that this approach effectively reduces the influence of extreme values while providing reliable error bars, resulting in more stable and robust estimates. We automate the whole procedure described above employing the adaptive KDE framework for density estimation of bootstrap resampled eigenvalues. In the case of Gaussian-convoluted histogram, the smearing parameter ($\sigma$) for Gaussian kernel in (Eq. 11 in III.1) needs to be provided manually.
V Results
Here, we present results obtained using our proposed method, applied to a wide range of two-point correlation functions, with the aim of extracting the corresponding energy levels of the associated states.
Before presenting the results obtained using the TGEVP method, it is worthwhile to first compare its effectiveness in extracting energy levels with that of other existing approaches. We begin by comparing results obtained from a given set of two-point correlation functions using different analysis methods: the standard effective mass method, Prony’s method [19, 21], Lanczos method, as described in Ref. [1], and the TGEVP method incorporating the proposed eigenvalue filtration. The effective mass is defined as $m_{eff}(t)={\mathrm{log}}[C(t)/C(t+1)]$, where $C(t)$ denotes a two-point correlation function. In the TGEVP method, eigenvalues are computed using two independent approaches—scipy.linalg.eig [41] and the FEAST eigensolver [42]—to cross-validate results and ensure high-precision evaluation. Subsequently, we discuss results from different correlation functions across a variety of hadrons, including nuclei.
V.1 Comparison between Prony’s method and TGEVP method
In this subsection, we present a comparative analysis of commonly used methods for extracting energy levels from lattice QCD two-point correlation functions: the traditional Prony method [19, 21] and the TGEVP framework introduced in this work. Our aim is to assess the relative reliability and stability of these approaches in determining ground-state energies.
The Prony method models the correlator as a sum of exponentials, reconstructing the parameters (masses and amplitudes) by solving a system of non-linear equations derived from the exponential time dependence of the two point correlation function. It is particularly effective at short time separations and allows for straightforward multi-state fits, but it is sensitive to statistical noise and the choice of fit window as is shown below.
To illustrate the comparison, we apply both methods to a same set of nucleon two-point correlation functions computed using clover valence quarks on HISQ gauge action. Figure 3 presents the results: the red point marks the TGEVP estimate, while the light green points indicate the Prony method’s estimates as a function of the minimum time slice $t_{\text{min}}$ included in the fit. At small $t_{\text{min}}$, the TGEVP estimates exhibit fluctuations due to excited-state contamination but quickly stabilize. It also agrees with the effective mass plateau within uncertainties, but is more stabilized. For both the TGEVP and Prony methods, the 16th and 84th percentiles are shown as error bars. Both the Prony and TGEVP estimates converge faster than the effective mass. However, the Prony method is found to be less stable than the TGEVP method, particularly at intermediate time slices. This instability becomes more pronounced at larger time separations, where the Prony results are consistently noisier.
Figure 3: Comparison of ground-state energy extraction using TGEVP (red), the two-state Prony method (blue) and effective mass (green) for a given set nucleon two-point functions. The $x$-axis corresponds to the minimum time slice $t_{\text{min}}$ for the Prony fits and the iteration index $m$ for the TGEVP estimate.
This comparison demonstrates that both the TGEVP and Prony methods can yield consistent estimates of the ground-state energy when applied appropriately. However, the TGEVP method exhibits greater robustness against excited-state contamination at early times and provides a cleaner signal at intermediate and large time separations compared to the Prony method.
V.2 Numerical comparison between Lanczos and TGEVP methods